OpenStudy (anonymous):

how do you solve 2^3/2?

OpenStudy (anonymous):

You have 2^3/2^1. So subtract the exponents giving: 2^(3-1)=2^2=4

OpenStudy (moonlight):

4

OpenStudy (moonlight):

2^3 = 8 divide by 2 4

OpenStudy (anonymous):

my teacher told me that the answer is \[\sqrt{8}\] so i was wondering how you get that but the answer isnt 4.